OpenStudy (love10129151):

Task 1 Each one is a different problem!? Solve using two different methods! Explain which method you found to be more efficient? a.3x-9y=3 6x-3y=--24 b. 7x-3y=20 5x+3y=16 c. y=1/2x-6 2x+6y=19

OpenStudy (texaschic101):

(a) elimination : 3x - 9y = 3 ---->(-2) 6x - 3y = -24 -------------- -6x + 18y = -6 (result of multiplying by -2) 6x - 3y = -24 -------------add 15y = - 30 y = - 2 sub in -2 for y 3x - 9y = 3 3x - 9(-2) = 3 3x + 18 = 3 3x = -18 + 3 3x = - 15 x = - 5 substitution 3x - 9y = 3 3x = 9y + 3 x = 9/3y + 3/3 x = 3y + 1 now sub 3y + 1 in for x 6x - 3y = -24 6(3y + 1) - 3y = -24 18y + 6 - 3y = -24 18y - 3y = -24 - 6 15y = -30 y = -2 now sub -2 in for y 3x - 9y = 3 3x - 9(-2) = 3 3x + 18 = 3 3x = 3 - 18 3x = -15 x = -5 I believe substitution is more efficient in this problem because when you isolate the variable, there is no fractions involved. Therefore, I find it easier to do. I am not going to do anymore of these....I think I left you a good example to go by. Good - luck

OpenStudy (mathstudent55):

@texaschic101 It looks like elimination is more efficient because it took fewer steps to solve the system of equations.

OpenStudy (texaschic101):

that might be because I included EVERY step. Some steps you could probably combine, making fewer steps.